Fresh / Soph / Jr / Sr / Grad (circle one) College and Major:

The grade I expect to receive is:

Mid-Semester Evaluation

36-617: Applied Linear Regression October 16, 2018

• This evaluation form has only one purpose: to improve the quality of instruction of this course, both for the rest of this semester, and for future semesters.

• Below are several general statements about the course. Please circle your response to each statement using the 5-point scale indicated.

• Use the blank space on the next page, to offer further comments & suggestions on the course.

• Your frank comments and constructive criticism would be greatly appreciated.

Do not sign your name.

Overall My understanding of the material covered to date:

1  2  3  4  5

I’m totally lost Challenging, but I think I’m understanding the material This course is really easy

My interest in, and enjoyment of, this course to date:

1  2  3  4  5

I’m bored stiff It’s OK but I like another course better One of my favorite classes

Course

1.

The course is not too hard and not too easy.

Disagree  Agree
Strongly  Strongly
1 2 3 4 5

If you disagree, is it too \underline{hard} or too \underline{easy}? .........................

2. The course is not too slow and not too fast.  1 2 3 4 5

If you disagree, is it too \underline{slow} or too \underline{fast}? .........................

3.

The things I am learning in this course will be useful in other classes or outside of school.

1 2 3 4 5 36-617: Applied Linear Regression Mid-Semester Evaluation

Please use this page to provide any other comments or suggestions on the course (positive or negative) that you think would be useful. You might think about the level of difficulty or easiness of the material, the instructor, the TA, the lectures, the textbooks, the homeworks, the projects, the classroom setup, or any other aspect of the course.